OpenStudy (anonymous):

What is the slope of the line between (-4, 4) and (-1, -2)?

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(-4\quad ,&4)\quad &(-1\quad ,&-2) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ y-y_1=m(x-x_1)\quad \textit{plug in your values and solve for "y"} \)

OpenStudy (anonymous):

\[slope = \frac {rise} {run}\] rise means the change in the y direction from one pint to the next point. from the point (-4, 4) to the point (-1, -2) there is a change in the y of -6 units. becuase we are going from 4 to -2, which is a difference of -6. so rise=-6 can u try to find the 'run' of the two points you have? or in other words find how many numbers are between -4 and -1

OpenStudy (anonymous):

these are my options 1 2 -2 -1

OpenStudy (jdoe0001):

hmm ohhh. it's just the slope, not the equation... hmm \(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(-4\quad ,&4)\quad &(-1\quad ,&-2) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

i think it would be 2?

OpenStudy (jdoe0001):

what did you get?